OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = -67?

OpenStudy (inkyvoyd):

an=a1+d(n-1)

OpenStudy (anonymous):

given a1=13 and a17=-67 thus 16d=-67-13=-80 hence d=-5 a35 =a1+34d =13+34(-5) =13-170 =-157 is ithe answer -157

OpenStudy (callisto):

Same type of question as the previous one. an = a1 + (n-1)d a17 = 13 + (17-1)d -67 = 13 + 16d Solve d. Then, apply the formula again a35 = a1 + (35-1)d = 13 + 34 d. You can solve it

OpenStudy (campbell_st):

use the formula to find d, the common difference \[T _{n} = a + (n - 1) d\] you have a = 13, n = 17 and T17 = -67 then when you have d... use the formula again with n = 35, a = 13

OpenStudy (anonymous):